OpenStudy (anonymous):

-5 x 1x 11 x 4 Solve the equation step by step. List each step along with the property you used during that step.

OpenStudy (kuoministers):

the x are times right? -5 * 1 * 11 * 4 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

-5 x 1 x 11 x 4

OpenStudy (kuoministers):

there isnt much working out you can show here.. you could simply put that in your calculator. but here is some working out lol -5 * 1 * 11 * 4 -5 * 44 -220

OpenStudy (kuoministers):

i dont think this is right cause u said this is an equation?

OpenStudy (kuoministers):

maybe can u draw the equation on paint and attach it as a file? lol

OpenStudy (anonymous):

How do you do that

OpenStudy (kuoministers):

so all the x's are all times not the pro-numeral X

OpenStudy (anonymous):

Ok something else (-4)(-4) write use exponents then simplify

OpenStudy (anonymous):

All times

OpenStudy (kuoministers):

The exponent of a number says how many times to use the number in a multiplication. In 8^2 the "2" says to use 8 twice in a multiplication, so 8^2 = 8 × 8 = 64 Exponents are also called Powers or Indices. ok (-4)(-4) is the same thing as (-4)^2 and -4 times -4 is 16

OpenStudy (anonymous):

Huh

OpenStudy (kuoministers):

u know what exponents is in the first place? its powers of a number So (-4)^2 the exponent is 2 then we do -4 times -4. Try that on ur calculator -4 * -4 we get 16.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So how do I simplify this that's when I get confused

OpenStudy (kuoministers):

another word for simplify is solve so just solve this equation.